OpenStudy (anonymous):

solve dy/dx=6x/(y^2(3-x^2)) for y(1)=1

hartnn (hartnn):

do you know variable separable method ? did you try to separate the variables for this DE ?

OpenStudy (anonymous):

yep, I got \[- \sqrt[3]{3 \ln (3-x ^{2})-3c}\]

OpenStudy (anonymous):

is that right?

hartnn (hartnn):

ok, now you can use the info : y(1) =1 which means , when x= 1, y = 1 you can plug these in the equation you got , to get the value of c

hartnn (hartnn):

let me check y^3/3 = -3 ln (3-x^2)+ c is that what you got first ?

OpenStudy (anonymous):

I got c= \[3\ln (2)+1/3\]

hartnn (hartnn):

you can keep the equation as y^3/3 = -3 ln (3-x^2)+ c to get the value of c

OpenStudy (anonymous):

yeah, that's what I got

hartnn (hartnn):

so 1/3 = -3 ln 2 +c yes, your 'c' is correct!

OpenStudy (anonymous):

really? i wasn't sure.. alrighty then! thank you!

hartnn (hartnn):

welcome ^_^ did you correctly got your final answer ?

OpenStudy (anonymous):

well, for the final answer, I got y = \[- (9 \ln (\frac{ 3-x ^{2} }{ 2 }) + 1)^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

yeah, not entirely sure for that bit

hartnn (hartnn):

let me reshuffle terms to see whether they can be simplified y^3/3 = -3 ln (3-x^2)+ 3 ln 2 +1/3 (y^3-1) = -9 [ln (3-x^2) -ln 2 ] y^3 -1 = -9 ln [(3-x^2)/2] y^3 = 1-9ln [(3-x^2)/2] \(\Large y = \sqrt[3]{1-9\ln (\dfrac{3-x^2}{2})}\)

hartnn (hartnn):

looks like you almost got the same thing, just be careful with the brackets

hartnn (hartnn):

\( (-9 \ln (\frac{ 3-x ^{2} }{ 2 }) + 1)^{\frac{ 1 }{ 3 }}\)

OpenStudy (anonymous):

ah yeah, I think that's why I got the second part wrong.. it asks to solve dy/dx - x^2y - xy = 0.. Thanks! :)

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

dy/dx = y (x^2-x) again variable seperable 1/y dy = (x^2-x) dx piece of cake from here :P

OpenStudy (anonymous):

sorry to bother again but I can't seem to solve the IVP \[dy/dx - x^{2}y - xy = 0\]

hartnn (hartnn):

add x^2y + xy on both sides first

OpenStudy (anonymous):

ahh ok, and solve it from there, yeah? i see it... i'll try it now :)

OpenStudy (anonymous):

y is the equation I found before yeah?

hartnn (hartnn):

i thought this one is completely new and different question....

OpenStudy (anonymous):

oh sorry, it's not.. it's a continuation of the previous one

hartnn (hartnn):

thats confusing...anyways, keep y as y only, as of now... factor out y from right side

OpenStudy (anonymous):

\[\frac{ 6x }{ y(3-x ^{2}) } = x ^{2} + x\] am I on the right track?

OpenStudy (anonymous):

hold on.. i think i'm wrong

OpenStudy (anonymous):

okay, so dy/dx = y(x^2 + x)

hartnn (hartnn):

you need to solve to find expression for 'y' again , in terms of x, right ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

ok, then separate the variables now

hartnn (hartnn):

dy/dx = y(x^2 + x) (1/y)dy = (x^2+x) dx and integrate both sides

OpenStudy (anonymous):

In(y) = (x^3)/3 + (x^2)/2

hartnn (hartnn):

correct. don't forget the +c

OpenStudy (anonymous):

oh crap.. yep, my bad. thanks

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

umm.. then how would I continue?

hartnn (hartnn):

\(\large \ln a = b \implies a = e^b \)

OpenStudy (anonymous):

alright..

hartnn (hartnn):

In(y) = (x^3)/3 + (x^2)/2 = x^2/6 (2x+3) + c \(\large y = c e^{[\dfrac{x^2}{6} (2x+3)]}\)

OpenStudy (anonymous):

\[y = e^{\frac{ x^3 }{ 3 } + \frac{ x^2 }{ 2 } + c}\]

hartnn (hartnn):

ok, then you can use \(e^{a+b} = e^a\times e^b\) so you'll get e^c, whic is also a constant, and you can write it as c only

hartnn (hartnn):

\(\Large y = e^{\frac{ x^3 }{ 3 } + \frac{ x^2 }{ 2 } + c} \implies y = e^{\frac{ x^3 }{ 3 } + \frac{ x^2 }{ 2 } }\times e^c \\ \Large y =c e^{\frac{ x^3 }{ 3 } + \frac{ x^2 }{ 2 }} \)

hartnn (hartnn):

if you want, you can simplify the exponent

OpenStudy (anonymous):

yep! I got that too :)

hartnn (hartnn):

if you simplify it, you'll get \(\large y = c e^{[\dfrac{x^2}{6} (2x+3)]}\)

OpenStudy (anonymous):

oh alright then

OpenStudy (anonymous):

and then would I use the the y(1) = 1 to solve for c?

hartnn (hartnn):

if y(1) = 1 is also given for this question, then yes

OpenStudy (anonymous):

c=e^(-5/6)

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

so \[y = e^\frac{ -5 }{ 6 } \times e^ (\frac{ x^2 }{ 6 } (2x+3))\]

hartnn (hartnn):

yeah, you can combine the exponents \(\large y = e^{[\dfrac{x^2}{6} (2x+3)]-\dfrac{5}{6}} =e^{\dfrac{x^2(2x+3)-5}6} \)

OpenStudy (anonymous):

ahh yep, got that :)